# Submitted toTopology Proceedings

### v2

@@ 3. Global elliptic sectors
By Definition 3.1, $\subset$" is clear. Since each of the $22m-$ 2 distinct global elliptic sectors lies between two definite directions and $m\geq 3$, there exists exactly one global elliptic sector $\mathcal{S}_{0}$ between $\theta_{+}$ and $\theta_{-}$, cf. Corollary 3.11. If there was a second one, there would be only $22<2m-$ 2 global elliptic sectors in $a$. Thus, we must have $\mathcal{S}(\theta_{+},\theta_{-},\Xi)=\mathcal{S}_{0}$. Moreover, by applying the second statement in Corollary 3.11, all $(\theta_{+},\theta_{-})$-elliptic orbits must lie in one of these $22m-$ 2 global elliptic sectors. Since the definite directions on the left hand side coincide with those on the right hand side in (3.2), all these orbits have to lie in $\mathcal{S}_{0}$ and thus in $\mathcal{S}(\theta_{+},\theta_{-},\Xi)$. This shows $\supset$". ∎
This second corollary shows that we need to choose a sector-forming orbit only in the case $m=2$. Here we have only two definite directions, which requires the choice of one arbitrary elliptic orbit to specify one of the two global elliptic sectors. The two sectors cannot be distinguished on the basis of the definite directions alone. If we consider the set on the right hand side of (3.2) in the case $m=2$, then this set is still open and unbounded, but not connected. The set would become connected (but no longer open) only by adding the equilibrium $a$. For this reason, the construction in Definition 3.1 uses a sector-forming orbit.
We observe that the boundaries of two adjacent global elliptic sectors do not have to coincide near $a$, as it is the case for the characteristic orbits in a FED, cf. [10, Theorem 4.4]. However, if they indeed coincide between all adjacent sectors, the local structure of the FED can even be transferred to the global phase portrait. The question therefore arises as to what the global geometric structure looks like in a region between two adjacent global elliptic sectors having no common boundary near $a$. How many unbounded orbits can exist in sum? This question takes up the discussions at the end of [10, Example 4.7] and [10, Example 4.8].
@@ 4. Basin of nodes and foci
This chapter provides detailed proofs of the topological results given in [4, Theorem 4.3]. In all claims and arguments of this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and a node or focus $a\in\mathbb{C}$ of (1.1).
@@
The stable (unstable) basin $\mathcal{N}$ of $F$ in $a$ is defined as
@@
The boundary of $a$ is defined as the boundary of the basin of $F$ in $a$.
@@
Assume w.l.o.g. that $a$ is stable. By [10, Definition 3.1], there exists $\varepsilon_{1}>0$ such that for all $y\in\mathcal{B}_{\varepsilon_{1}}(a)$ we have $w_{+}(\Gamma(y))=\{a\}$. This proves $\mathcal{B}_{\varepsilon_{1}}(a)\subset\mathcal{N}$ and $a\in\mathring{\mathcal{N}}$. Let $x\in\mathcal{N}\setminus\{a\}$ be arbitrary. Then there exists a $\tau>0$ with $\xi:=\Phi(\tau,x)\in\mathcal{B}_{\varepsilon_{1}}(a)$. Choose $\varepsilon_{2}:=\frac{1}{2}\min\{|\xi-a|,\varepsilon_{1}-|\xi-a|\}>0$ such that $\mathcal{B}_{\varepsilon_{2}}(\xi)\subset\mathcal{B}_{\varepsilon_{1}}(a)$. By continuity of the flow, cf. [19, Chapter 2.4, Theorem 4], there exists a $\delta>0$ such that $|\Phi(\tau,z)-\xi|<\varepsilon_{2}$ for all $z\in\mathcal{B}_{\delta}(x)$. Thus $w_{+}(\Gamma(z))=\{a\}$ for all $z\in\mathcal{B}_{\delta}(x)$ and $\mathcal{N}$ is open.
For $x\in\mathcal{N}$ set $J_{x}:=\Gamma(x)\cup\{a\}$ and define the cover $C:=\{J_{x}:x\in\mathcal{N}\}$ of $\mathcal{N}$. Since $a$ lies on the boundary of each orbit in $\mathcal{N}$, every set in $C$ is connected, cf. [18, Theorem 23.4]. Moreover, for $J,\tilde{J}\in C$ it always holds $J\cap\tilde{J}\supset\{a\}\not=\emptyset$. Thus, by [18, Theorem 23.3], $\mathcal{N}$ is connected. From [23, Proposition 12.25] it follows that $\mathcal{N}$ is even path-connected. ∎
Lemma 4.4.
All centers $\tilde{a}\in F^{-1}(\{0\})\setminus\{a\}$ with its corresponding basin $\mathcal{V}$ satisfy $\mathcal{V}\cap\partial\mathcal{N}=\emptyset$. In particular, there are no centers on $\partial\mathcal{N}$.
Proof.
Assume w.l.o.g. that $a$ is stable. Suppose there exists a center $\tilde{a}\in\mathbb{C}{}\setminus\{a\}$ such that its basin $\mathcal{V}$ satisfies $\mathcal{V}\cap\partial\mathcal{N}\not=\emptyset$. Since $\mathcal{V}$ is open, cf. Theorem 2.4, we can choose $z\in\mathcal{V}\cap\partial\mathcal{N}$ and $r>0$ sufficiently small such that $\mathcal{B}_{r}(z)\subset\mathcal{V}$. Since $z\in\partial\mathcal{N}$, there exists $y\in\mathcal{N}\cap\mathcal{B}_{r}(z)\not=\emptyset$ satisfying $w_{+}(\Gamma(y))=\{a\}$. But this is impossible for a periodic orbit in $\mathcal{V}$. Thus $\mathcal{V}\cap\partial\mathcal{N}=\emptyset$. In particular, there are no centers on $\partial\mathcal{N}$, i.e. all equilibria on the boundary of $a$ have to be nodes, foci or zeros of order $m\geq 2$. ∎
In the proof of [4, Theorem 4.3] the author does not investigate the case of an isolated equilibrium on the boundary of a node or focus. This is a significant gap, which is closed by the following result.
@@
$\forall\,\tilde{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\}):\forall\,\rho>0:\left(\mathcal{B}_{\rho}(\tilde{a})\cap\partial\mathcal{N}\right)\setminus\{\tilde{a}\}\not=\emptyset,$
@@
$\partial\mathcal{N}\cap(\underbrace{\mathcal{B}_{\rho}(\tilde{a})\setminus\{\tilde{a}\}}_{=:A})=\left(\mathcal{B}_{\rho}(\tilde{a})\cap\partial\mathcal{N}\right)\setminus\{\tilde{a}\}=\emptyset\text{.}$
It follows [^6] The simple but technical proof of ^$\hat{\partial}(\mathcal{N}\cap A)\subset\partial\mathcal{N}\cap A=\emptyset$, where $\hat{\partial}$ denotes the boundary with respect to the subspace topology on $A$, is omitted. that $\mathcal{N}\cap A\not=\emptyset$ is open and closed with respect to the subspace topology on $A$. Since the punctured circle $A$ is connected, we must have $\mathcal{N}\cap A=A$. Hence, since $\tilde{a}\not\in\mathcal{N}$, we conclude
(4.1) $\mathcal{N}\cap\mathcal{B}_{\rho}(\tilde{a})=\mathcal{B}_{\rho}(\tilde{a})\setminus\{\tilde{a}\}\text{.}$
If $\tilde{a}$ was of order $m\geq 2$, then, by [10, Theorem 4.4], there would be at least one homoclinic orbit in $A$ tending to $\tilde{a}$ in both time directions, i.e. equation (4.1) would not be fulfilled and $\tilde{a}$ would not be isolated. Thus, $\tilde{a}$ must be a node or focus, cf. Lemma 4.4. Denote by $\tilde{\mathcal{N}}$ the basin of $F$ in $\tilde{a}$. We reduce $\rho$ such that $\mathcal{B}_{\rho}(\tilde{a})\subset\tilde{\mathcal{N}}$ and $\mathcal{B}_{\rho}(a)\subset\mathcal{N}$. By (4.1) and Proposition 4.2, we have $\tilde{\mathcal{N}}\subset\mathcal{N}$.
By applying the theory of circles without contact around nodes and foci, cf. [2, §3, 10.-14., §7, 1.-2. and §18, Lemma 3], we find two continuously differentiable closed paths $C_{1}\subset\mathcal{B}_{\rho}(a)$ and $C_{2}\subset\mathcal{B}_{\rho}(\tilde{a})$, $C_{1}\cap C_{2}=\emptyset$, being nowhere tangential to $F$ and satisfying $\operatorname{Int}(C_{1})\cap F^{-1}(\{0\})=\{a\}$ and $\operatorname{Int}(C_{2})\cap F^{-1}(\{0\})=\{\tilde{a}\}$. [^7] In particular, from the equations (6) and (11) in [2, §7, 1.] and the remarks made in [2, §7, 2.] it follows that $C_{1}$ as well as $C_{2}$ can be chosen as a linear transformed circle or ellipse. It holds $\overline{\operatorname{Int}(C_{1})}\cap\overline{\operatorname{Int}(C_{2})}=\emptyset$. Moreover, every orbit in $\mathcal{N}$ ($\tilde{\mathcal{N}}$) crosses $C_{1}$ ($C_{2}$) exactly once, cf. [2, §3, 10., Figure 54]. We equip $C_{1}$ and $C_{2}$ with the subspace topology. Using this construction, the map $\Psi:\mathcal{N}\setminus\{a\}\to\mathbb{R}$ is well-defined, when $\Psi(x)$ is the transit time888cf. [4, Definition 3.2]. the point $x\in\mathcal{N}\setminus\{a\}$ needs to $C_{1}$ along the path $\Gamma(x)$, i.e. $\{\Phi(\Psi(x),x)\}=\Gamma(x)\cap C_{1}$. Moreover, we define $\eta:C_{2}\to C_{1}$ by $\eta(x):=\Phi(\Psi(x),x)$. By the continuous dependence of the flow on initial conditions, cf. [19, Chapter 2.4, Theorem 4], and the continuous dependence of the transit time integral on the integral limits, $\Psi$ and $\eta$ are both continuous.
One the one hand, since $C_{2}$ is compact, also $\eta(C_{2})$ is compact in $C_{1}$, cf. [18, Theorem 26.5]. Thus, $\eta(C_{2})$ is closed in $C_{1}$, cf. [18, Theorem 26.3]. On the other hand, by Proposition 4.2, we have $\eta(C_{2})\subset\tilde{\mathcal{N}}\cap C_{1}$. Additionally, since all orbits through $\tilde{\mathcal{N}}\cap C_{1}\subset\tilde{\mathcal{N}}$ have to cross $C_{2}$, we also have $\tilde{\mathcal{N}}\cap C_{1}\subset\eta(C_{2})$, i.e. $\eta(C_{2})=\tilde{\mathcal{N}}\cap C_{1}$. Moreover, by Theorem 4.3, $\tilde{\mathcal{N}}\cap C_{1}$ is open in $C_{1}$. We conclude that $\eta(C_{2})\not=\emptyset$ is open and closed in $C_{1}$. Since $C_{1}$ is connected, we must have $C_{1}=\eta(C_{2})$, i.e. $\eta$ is surjective. Let $y_{1},y_{2}\in C_{2}$ be arbitrary. Assume $\eta(y_{1})=\eta(y_{2})$. Then, by definition of $\eta$, $\Gamma(y_{1})=\Gamma(\eta(y_{1}))=\Gamma(\eta(y_{2}))=\Gamma(y_{2})$. Since all orbits in $\tilde{\mathcal{N}}$ cross $C_{2}$ exactly once, we must have $y_{1}=y_{2}$. Hence, $\eta$ is also injective. It follows that $\eta$ is bijective and thus a homeomorphism, cf. [18, Theorem 26.6]. We conclude that every orbit in $\mathcal{N}$ and $\tilde{\mathcal{N}}$ crosses $C_{1}$ as well as $C_{2}$ exactly once, i.e. [^9] The set $\mathcal{H}$ is the so-called heteroclinic region between $a$ and $\tilde{a}$ and is introduced in Chapter 5.
(4.2) $\mathcal{H}:=\mathcal{N}\cap\tilde{\mathcal{N}}=\mathcal{N}\setminus\{a\}=\tilde{\mathcal{N}}\setminus\{\tilde{a}\}=\bigcup\limits_{\mathclap{x_{2}\in C_{2}}}\Gamma(x_{2})=\bigcup\limits_{\mathclap{x_{1}\in C_{1}}}\Gamma(x_{1})\text{.}$
Since $C_{1}\subset\mathcal{H}$ is a Jordan curve, there exists a homeomorphism $h:C_{1}\to S^{1}$. Equip the cylinder surface $B:=\mathbb{R}\times S^{1}\subset\mathbb{R}^{3}$ with the subspace topology and define the map $\zeta:B\to\mathcal{H}$ by $\zeta(t,x):=\Phi(t,h^{-1}(x))$. By (4.2), we have $\zeta(B)=\mathcal{H}$, i.e. $\zeta$ is surjective. Let $(t_{1},x_{1}),(t_{2},x_{2})\in B$ be arbitrary and assume $\zeta(t_{1},x_{1})=\zeta(t_{2},x_{2})$. Then $h^{-1}(x_{1}),h^{-1}(x_{2})\in C_{1}$ and
Since all orbits in $\mathcal{H}$ cross $C_{1}$ exactly once, we get $h^{-1}(x_{1})=h^{-1}(x_{2})$. Since $h$ is bijective, we conclude $x_{1}=x_{2}$. This gives us $\Phi(t_{1},h^{-1}(x_{1}))=\Phi(t_{2},h^{-1}(x_{1}))$ and thus $t_{1}=t_{2}$, since $\Gamma(h^{-1}(x_{1}))$ is not periodic. It follows that $\zeta$ is also injective and thus bijective. Hence, there exists the inverse $\zeta^{-1}$ of $\zeta$, which is given by $\zeta^{-1}(x):=(\Psi(x),h(\Phi(\Psi(x),x)))$, $x\in\mathcal{H}$. As a composition of continuous functions, $\zeta$ as well as $\zeta^{-1}$ are continuous and therefore $\zeta$ is a homeomorphism.
By [18, Corollary 52.5], the fundamental group of $\mathcal{H}$ is isomorphic to $B$. Additionally, as the circle line $S^{1}\subset\mathbb{R}^{2}$ is a deformation retract of the cylinder surface $B$, we conclude that the fundamental group of $\mathcal{H}$ is isomorphic to the additive group $\mathbb{Z}$, cf. [18, Theorem 54.5] and [18, Theorem 58.3]. But $\mathcal{H}$ has the two holes $a$ and $\tilde{a}$, which is impossible for a set having a fundamental group isomorphic to $\mathbb{Z}$. In particular, by using (4.2), we can choose the loop $C_{2}\cup\Phi([0,\Psi(z)],z)\cup C_{1}\subset\mathcal{H}$ with an arbitrary point $z\in C_{2}$. This loop has $a\not\in\mathcal{H}$ as well as $\tilde{a}\not\in\mathcal{H}$ in its interior and thus cannot be transferred to a loop on $S^{1}$. All in all, we derive a contradiction and $\tilde{a}$ cannot be isolated on $\partial\mathcal{N}$.
Moreover, all equilibria $\hat{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$ are nodes, foci or equilibria of order $m\geq 2$, cf. Lemma 4.4. Thus, we can choose $\hat{\rho}$ such that for all $\xi\in\mathcal{B}_{\hat{\rho}}(\hat{a})$ there holds $\hat{a}\in w_{+}(\Gamma(\xi))\cup w_{-}(\Gamma(\xi))$, cf. [10, Definition 3.1] and [10, Theorem 4.4]. If we choose $\xi\in(\mathcal{B}_{\hat{\rho}}(\hat{a})\cap\partial\mathcal{N})\setminus\{\hat{a}\}\not=\emptyset$, then $\Gamma:=\Gamma(\xi)\subset\mathcal{N}$ satisfies $\hat{a}\in w_{+}(\Gamma)\cup w_{-}(\Gamma)$, i.e. $\hat{a}$ is attached to $\Gamma$. ∎
@@
If $\partial\mathcal{N}=\emptyset$, then $\mathcal{N}$ is open, closed and nonempty, i.e. $\mathcal{N}=\mathbb{C}{}$ is clearly unbounded. Thus, assume $\partial\mathcal{N}\not=\emptyset$. Additionally, assume w.l.o.g. that $a$ is stable.
Suppose there exists a bounded orbit $\Gamma\subset\partial\mathcal{N}\setminus F^{-1}(\{0\})$. As in the proof of Theorem 2.6, $K:=\overline{\Gamma}$ is compact and $\Gamma\cup w_{\pm}(\Gamma)\subset K\subset\partial\mathcal{N}$. By Lemma 4.4, $\Gamma$ cannot be periodic. Hence, by [10, Theorem 5.1], there exist two equilibria $a_{+},a_{-}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$ such that $w_{\pm}(\Gamma)=\{a_{\pm}\}$, i.e. $\Gamma$ is either heteroclinic or homoclinic. In both cases we want to derive a contradiction.
Suppose $\Gamma$ is homoclinic, i.e. $a_{+}=a_{-}$. By Corollary 3.11, $\Gamma$ is a sector-forming orbit lying in an global elliptic sector $\mathcal{S}_{0}$ in $a_{+}$. Choose $z\in\Gamma\subset\mathcal{S}_{0}\cap\partial\mathcal{N}$. Since $\mathcal{S}_{0}$ is open, cf. Theorem 3.6, there exists $\varepsilon>0$ such that $\mathcal{B}_{\varepsilon}(z)\subset\mathcal{S}_{0}$. Thus, analogous to the proof of Lemma 4.4, we can choose a point $y\in\mathcal{S}_{0}\cap\mathcal{N}\not=\emptyset$, i.e. $\Gamma(y)\subset\mathcal{N}$ is a homoclinic orbit. This is a contradiction.
Suppose, $\Gamma$ is heteroclinic, i.e. $a_{+}\not=a_{-}$. Since $a_{+},a_{-}\in\partial\mathcal{N}$ and $a\in\mathring{\mathcal{N}}$, cf. Theorem 4.3, it holds $a\not\in\{a_{+},a_{-}\}$. Thus, $a_{+}$ cannot be a node or focus, since its basin would have nonempty intersection with $\mathcal{N}$ [^10] If we investigated the case that $a$ is unstable, we would choose $a_{-}$ instead of $a_{+}$ at this point. Note that an orbit in $\mathcal{N}$ cannot have $a$ as well as $a_{+}$ in its limit set for $t\to\infty$. Additionally, by Lemma 4.4, $a_{+}\in\partial\mathcal{N}$ cannot be a center. It follows that $a_{+}$ must have order $m\geq 2$. By [10, Proposition 4.3], there exist $\theta\in\mathcal{E}(F,m)$ and $r,\delta>0$ such that $\Gamma$ as well as the orbit $\Gamma(x_{0})$ through $x_{0}\in\{x\in\mathbb{R}^{2}:|x-a_{+}|<r,|\arg(x-a_{+})-\theta|<\delta\}=:A$ tends to $a_{+}$ for $t\to\infty$. Choose $\hat{z}\in\Gamma\cap A\subset\partial\mathcal{N}\cap A$. Since $A$ is open, we can choose $\hat{\varepsilon}>0$ such that $\mathcal{B}_{\hat{\varepsilon}}(\hat{z})\subset A$. Thus, again analogous to the proof of Lemma 4.4, we can choose $x_{0}\in A\cap\partial\mathcal{N}\not=\emptyset$, i.e. $\Gamma(x_{0})\subset\mathcal{N}$ tends to $a_{+}\not=a$ for $t\to\infty$. This is a contradiction.
We conclude that $\Gamma$ cannot be bounded. All in all, it follows that $\partial\mathcal{N}$ consists only of equilibria (but no centers) and unbounded orbits. Since $\partial\mathcal{N}\not=\emptyset$, we can apply Proposition 4.5 to conclude that there exists at least one unbounded orbit lying on the boundary of $a$. Thus, also $\mathcal{N}$ is unbounded. ∎
@@
Let $L\subset\mathcal{N}$ be a loop. Since $L$ is compact and $\mathcal{N}$ is open, cf. Theorem 4.3, there exist points $y_{1},\ldots,y_{\ell}\in L$ and radii $r_{1},\ldots,r_{\ell}>0$, $\ell\in\mathbb{N}{}$, such that
$L\subset\bigcup_{k=1}^{\ell}\mathcal{B}_{r_{k}}(y_{k})=:O$
and $\overline{O}\subset\mathcal{N}$. The latter property can be ensured by reducing the radii. In addition, there exists a closed Jordan curve $J\subset\partial O\subset\mathcal{N}$ such that $L\subset\operatorname{Int}(J)$. If $\operatorname{Int}(J)\setminus\mathcal{N}\not=\emptyset$, we can choose $x\in\partial\mathcal{N}\cap\operatorname{Int}(J)\not=\emptyset$. [^11] A detailed topological argumentation for this can be found in the proof of [9, Theorem 5.10]. If $x$ is not an equilibrium, $\Gamma(x)$ must be unbounded, cf. Theorem 4.6. If $x$ is an equilibrium, there exists an orbit $\Gamma\subset\partial\mathcal{N}$ with $x$ in its limit set, cf. Proposition 4.5. Thus, by openness of $\operatorname{Int}(J)$, we have $\Gamma\cap\operatorname{Int}(J)\not=\emptyset$. In both cases we get an unbounded orbit with nonempty intersection with $\operatorname{Int}(J)$. But $\operatorname{Int}(J)$ is bounded, i.e. this is a contradiction. Thus, we have $\operatorname{Int}(J)\subset\mathcal{N}$. Since $\operatorname{Int}(J)$ is simply connected, $L$ can be continuously deformed to a point. Since $L$ is arbitrary and $\mathcal{N}$ is already path-connected, it follows that $\mathcal{N}$ is even simply connected. ∎

### v3

@@ 3.3. Final Corollaries
The key point underlying this result is that, for $m\geq 3$, there exist more than two definite directions. Since each of the 2m22m-2 distinct unbounded global elliptic sectors lies between two adjacent definite directions, there always exists exactly one global elliptic sector $\mathcal{S}_{0}$ between two adjacent arbitrarily chosen angles $\theta_{+}$ and $\theta_{-}$, cf. Corollary 3.11. Moreover, by applying the second statement in Corollary 3.11, all $(\theta_{+},\theta_{-})$-elliptic orbits must lie in one of these 2m22m-2 global elliptic sectors. This shows (3.3) and (3.4). ∎
This second corollary shows that we need to choose a sector-forming orbit only in the case $m=2$. Here we have only two definite directions, which requires the choice of one arbitrary homoclinic orbit to specify one of the two global elliptic sectors. The two sectors cannot be distinguished on the basis of the definite directions alone. If we consider the set on the right hand side of (3.2) in the case $m=2$, then this set is still open and unbounded, but not connected. The set would become connected (but no longer open) only by adding the equilibrium $a$. For this reason, the construction in Definition 3.1 is based on the notion of a sector-forming orbit, which is defined independently of the choice of definite directions.
We observe that the boundaries of two adjacent global elliptic sectors in general do not have to coincide near $a$, as it is the case for the characteristic orbits in a FED. However, if they indeed coincide between all adjacent sectors, the local structure of the FED can even be transferred to the global phase portrait. The question therefore arises as to what the global geometric structure looks like in a region between two adjacent global elliptic sectors having no common boundary near $a$. How many separatrices can tend to a multiple equilibrium (either in positive or negative time)? This question takes up the discussions at the end of [19, Example 4.7] and [19, Example 4.8]. We will discuss these issues in Chapter 5.
Finally, we conclude by addressing the question of how the global elliptic sector fits into the theory of canonical regions.
Corollary 3.14.
Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. $\mathcal{S}$ is a canonical region of strip type. There exists a global smooth transversal through $\mathcal{S}$, i.e. an open smooth arc crossing each orbit in $\mathcal{S}$ exactly once.
Proof.
By openness of $\mathcal{S}$, all orbits in $\mathcal{S}$ are part of a strip canonical region $R$. Thus, there are no separatrices in $\mathcal{S}$. Since $\partial\mathcal{S}$ does not have any homoclinic orbits, we have $\partial\mathcal{S}\cap R=\emptyset$, i.e. $R=\mathcal{S}$. The rest follows directly from [9, Theorem 5.1] and [10, Theorem I]. ∎
@@ 4. Basin of nodes and foci
In this chapter, we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and a node or focus $a\in\mathbb{C}$ of (1.1). We provide detailed proofs of the topological results given in [5, Theorem 4.3].
@@
The stable (unstable) basin of attraction (repulsion) $\mathcal{N}$ of $F$ in $a$ is defined as
@@
The boundary of $a$ is defined as the boundary of $\mathcal{N}$.
@@
Assume w.l.o.g. that $a$ is stable. By [19, Definition 3.1], there exists $\varepsilon_{1}>0$ such that for all $y\in\mathcal{B}_{\varepsilon_{1}}(a)$ we have $w_{+}(\Gamma(y))=\{a\}$. This proves $\mathcal{B}_{\varepsilon_{1}}(a)\subset\mathcal{N}$ and $a\in\mathring{\mathcal{N}}$. Let $x\in\mathcal{N}\setminus\{a\}$ be arbitrary. Then there exists a $\tau>0$ with $\xi:=\Phi(\tau,x)\in\mathcal{B}_{\varepsilon_{1}}(a)$. Choose $\varepsilon_{2}:=\frac{1}{2}\min\{|\xi-a|,\varepsilon_{1}-|\xi-a|\}>0$ such that $\mathcal{B}_{\varepsilon_{2}}(\xi)\subset\mathcal{B}_{\varepsilon_{1}}(a)$. By continuity of the flow, cf. [29, Chapter 2.4, Theorem 4], there exists a $\delta>0$ such that $|\Phi(\tau,z)-\xi|<\varepsilon_{2}$ for all $z\in\mathcal{B}_{\delta}(x)$. Thus $w_{+}(\Gamma(z))=\{a\}$ for all $z\in\mathcal{B}_{\delta}(x)$ and $\mathcal{N}$ is open.
For $x\in\mathcal{N}$ set $J_{x}:=\Gamma(x)\cup\{a\}$ and define the cover $C:=\{J_{x}:x\in\mathcal{N}\}$ of $\mathcal{N}$. Since $a$ lies on the boundary of each orbit in $\mathcal{N}$, every set in $C$ is connected, cf. [28, Theorem 23.4]. Moreover, for $J,\tilde{J}\in C$ it always holds $J\cap\tilde{J}\supset\{a\}\not$$\emptyset$. Thus, by [28, Theorem 23.3], $\mathcal{N}$ is connected. From [33, Proposition 12.25] it follows that $\mathcal{N}$ is even path-connected. ∎
In the proof of [5, Theorem 4.3] the author does not investigate the case of an isolated equilibrium on the boundary of a node or focus. This is a significant gap, which is closed by the following result.
@@
$\forall\,\tilde{a}\in\partial\mathcal{N}\cap F^{-1}(\{0\}):\forall\,\rho>0:\left(\mathcal{B}_{\rho}(\tilde{a})\cap\partial\mathcal{N}\right)\setminus\{\tilde{a}\}\not$$\emptyset$,
@@
This proof, carried out by contradiction, relies on constructing a homeomorphism between two appropriately chosen circles without contact, each encircling a focus/node. Since this is highly technical, a very detailed step-by-step proof can be found in the appendix. ∎
@@
If $\partial\mathcal{N}=\emptyset$, then $\mathcal{N}$ is open, closed and nonempty, i.e. $\mathcal{N}=\mathbb{C}{}$ is clearly unbounded. Thus, assume $\partial\mathcal{N}\not$$\emptyset$. Additionally, assume w.l.o.g. that $a$ is stable.
Suppose there exists a bounded orbit $\Gamma\subset\partial\mathcal{N}\setminus F^{-1}(\{0\})$. As in the proof of Theorem 2.4, $K:=\overline{\Gamma}$ is compact and $\Gamma\cup w_{\pm}(\Gamma)\subset K\subset\partial\mathcal{N}$. By Lemma 2.6, $\Gamma$ cannot be periodic. Hence, by Theorem 1.1, there exist two equilibria $a_{+},a_{-}\in\partial\mathcal{N}\cap F^{-1}(\{0\})$ such that $w_{\pm}(\Gamma)=\{a_{\pm}\}$, i.e. $\Gamma$ is either heteroclinic or homoclinic. In both cases we want to derive a contradiction.
Suppose $\Gamma$ is homoclinic, i.e. $a_{+}=a_{-}$. By Corollary 3.11, $\Gamma$ is a sector-forming orbit lying in an global elliptic sector $\mathcal{S}_{0}$ in $a_{+}$. Choose $z\in\Gamma\subset\mathcal{S}_{0}\cap\partial\mathcal{N}$. By openness of $\mathcal{S}_{0}$, cf. Theorem 3.6, we get $\mathcal{S}_{0}\cap\mathcal{N}\not$$\emptyset$, contradicting the fact that all orbits in $\mathcal{N}$ tend to $a\not\in\{a_{+},a_{-}\}$.
Suppose, $\Gamma$ is heteroclinic, i.e. $a_{+}\not=a_{-}$. Since $\not\in\{a_{+},a_{-}\}, a+a_{+}$ cannot be a node or focus, since its basin would have nonempty intersection with $\mathcal{N}$. [^6] If we investigated the case that $a$ is unstable, we would choose $a_{-}$ instead of $a_{+}$ at this point. Additionally, by Lemma 2.6, $a_{+}\in\partial\mathcal{N}$ cannot be a center. It follows that $a_{+}$ must have order $m\geq 2$. By [19, Proposition 4.3], there exist $\theta\in\mathcal{E}(F,m)$ and $r,\delta>0$ such that $\Gamma$ as well as the orbit $\Gamma(x_{0})$ through $x_{0}\in\{x\in\mathbb{R}^{2}:|x-a_{+}|<r,|\arg(x-a_{+})-\theta|<\delta\}=:A$ tend to $a_{+}$ for $t\to\infty$. We have $\Gamma\cap A\not$$\emptyset$. Since $A$ is open, we can choose $x_{0}\in A\cap\mathcal{N}\not$$\emptyset$, i.e. $\Gamma(x_{0})\subset\mathcal{N}$ tends to $a_{+}\not=a$ for $t\to\infty$. This is a contradiction.
We conclude that $\Gamma$ cannot be bounded. All in all, it follows that $\partial\mathcal{N}$ consists only of equilibria (but no centers) and unbounded orbits. Since $\partial\mathcal{N}\not$$\emptyset$, we can apply Proposition 4.4 to conclude that there exists at least one unbounded orbit lying on the boundary of $a$. Thus, also $\mathcal{N}$ is unbounded. ∎
@@
Let $J\subset\mathcal{N}$ be a loop and suppose that $\operatorname{Int}(J)\setminus\mathcal{N}\not$$\emptyset$. Then we can choose $x\in\partial\mathcal{N}\cap\operatorname{Int}(J)\not$$\emptyset$. If $x$ is not an equilibrium, $\Gamma(x)$ must be unbounded, cf. Theorem 4.5. If $x$ is an equilibrium, there exists an orbit $\Gamma\subset\partial\mathcal{N}$ with $x$ in its limit sets, cf. Proposition 4.4. In both cases we get an unbounded orbit on the boundary of $a$ having nonempty intersection with $\operatorname{Int}(J)$. Since $\operatorname{Int}(J)$ is bounded, this orbit must intersect $J$. This is a contradiction to the openness of $\mathcal{N}$. Thus, we conclude $\operatorname{Int}(J)\subset\mathcal{N}$ and $J$ can be continuously deformed to a point in $\mathcal{N}$. Since $J$ is arbitrary and $\mathcal{N}$ is already path-connected, it follows that $\mathcal{N}$ is even simply connected. ∎
To put the basin $\mathcal{N}$ into the framework of the theory of canonical regions and global transversals, we consider the following example of a polynomial vector field.
Example 4.7.
We consider the entire vector field $F:\mathbb{C}{}\to\mathbb{C}{}$ with
$F(x):=x^{3}-x=x(x-1)(x+1)\text{.}$
Figure 2. Local phase portrait of system (1.1) with $F(x)=x^{3}-x$, plotted with Wolfram Mathematica.
All equilibria $F^{-1}(\{0\})=\{0,1,-1\}$ are nodes. The nodes at $\pm 1$ are repelling, while the node at $a=0$ is attracting. By using the notation $\mathrm{i}\mathbb{R}:=\{x\in\mathbb{C}:\Re(x)=0\}$, we obtain the following two properties of the phase space by a straightforward calculation:
  • $F(\alpha)\in\mathbb{R}$ for all $\alpha\in\mathbb{R}$, i.e. $\mathbb{R}$ is flow-invariant.
  • $F(\mathrm{i}\beta)\in\mathrm{i}\mathbb{R}$ for all $\beta\in\mathbb{R}$, i.e. $\mathrm{i}\mathbb{R}$ is flow-invariant.
Thus, (1,0)(-1,0) as well as $(0,1)$ are orbits with distinct negative limit sets ($\{1\}$ and $\{-1\}$, respectively). Moreover, the upper (denoted $A_{1}$) and lower (denoted $A_{2}$) parts of the imaginary axis are orbits, both having empty negative limit sets. All four of these orbits lie in the basin of the node at $a=0$. We therefore conclude that these orbits cannot belong to the same canonical region, cf. [9, Chapter II.3]. In particular, since the orbits through arbitrarily close neighborhoods of $A_{1}$ and $A_{2}$ have nonempty negative limit sets, $A_{1}$ and $A_{2}$ must be separatrices in the sense of [9]. A similar observation applies to foci if we consider the rotated vector field $\hat{F}(x):=(1+\mathrm{i})F(x)$, $x\in\mathbb{C}$.
Example 4.7 demonstrates that the basins of nodes and foci do not necessarily have to form canonical regions. Applying Proposition 4.4, we obtain the following result: if $\mathcal{N}$ is a canonical region, then there cannot be any equilibria on $\partial\mathcal{N}$.
In general, nodal and spiral canonical regions can be constructed with the simple linear vector field $x\mapsto\lambda x$ with $\Re(\lambda)\not=0$, cf. [19, Definition 3.1].